A charge Q is placed at the centre of a square. If electric field intensity due to the charge at the corners of
the square is E1 and the intensity at the mid point of the side of square is E2, then the ratio of E1/E2 will be

Grade:12

1 Answers

Arun
5 years ago
The distance between the charge and the corner is a√2
distance between midpoint of side and charge is a/2
so net field is given by
kQ/r^2

e1=kQ/(1/2)a^2
e2=kQ/(1/4)a^2
e1/e2= 2